OpenStudy (anonymous):

A family has two cars. The first car has a fuel efficiency of 20 miles per gallon of gas and the second has a fuel efficiency of 25 miles per gallon of gas. During one particular week, the two cars went a combined total of 1325 miles, for a total gas consumption of 60 gallons. How many gallons were consumed by each of the two cars that week?

OpenStudy (anonymous):

Example: A family has two cars. The first car has a fuel efficiency of 25 miles per gallon of gas and the second has a fuel efficiency of 20 miles per gallon of gas. During one particular week, the two cars went a combined total of 1300 miles, for a total gas consumption of 55 gallons. How many gallons were consumed by each of the two cars that week? gallons were consumed by each of the two cars that week are 40 and 15 Solution:Set the miles of the first car is X, the gallons consumed is Y then (55-Y)x 20 =1300- X Y x 25 = X so Y = 40, X = 1000

OpenStudy (anonymous):

hope it helps

OpenStudy (anonymous):

it does, thank you!

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

best answer please :)

OpenStudy (anonymous):

thanks